# GLiDRE: Generalist Lightweight model for Document-level Relation Extraction

### v1

@@ Abstract
Relation Extraction (RE) is a fundamental task in Natural Language Processing, and its document-level variant poses significant challenges, due to the need to model complex interactions between entities across sentences. Current approaches, largely based on the ATLOP architecture, are commonly evaluated on benchmarks like DocRED and Re-DocRED. However, their performance in zero-shot or few-shot settings remains largely underexplored due to the task’s complexity. Recently, the GLiNER model has shown that a compact NER model can outperform much larger Large Language Models. With a similar motivation, we introduce GLiDRE, a new model for document-level relation extraction that builds on the key ideas of GliNER. We benchmark GLiDRE against state-of-the-art models across various data settings on the Re-DocRED dataset. Our results demonstrate that GLiDRE achieves state-of-the-art performance in few-shot scenarios. Our code is publicly available [^1] https://github.com/robinarmingaud/glidre.
@@ 1 Introduction
DocRE presents a more realistic and complex scenario, especially relevant for real-world applications such as Information Extraction (IE) in biomedical or industrial texts, where relationships often span sentence boundaries and the entity pairs of interest are not pre-identified. The complexity of DocRE is due to the quadratic growth of negative pairs with the number of entities in a document. Evaluation is commonly conducted on datasets like DocRED (Yao et al., 2019) and its revised version Re-DocRED (Tan et al., 2022b), which reduces the issue of false negatives. These two datasets contain more complex documents with an average of 19 entities and approximately 390 candidate pairs per text, for a total of 3,053 annotated Wikipedia articles in the training set.
State-of-the-art approaches to DocRE generally extend the model ATLOP (Zhou et al., 2021), which introduces Localized Context Pooling, using attention mechanisms to enhance relation representations, and an adaptive-thresholding loss that adjusts decision thresholds dynamically based on entity pairs. Subsequent work has extended ATLOP in various directions. For example, DREEAM (Ma et al., 2023) uses evidences to enhance performance, while KD-DocRE (Tan et al., 2022a) uses axial attention to model interdependencies among entity pairs, adopts an adaptive focal loss to address class imbalance, and integrates knowledge distillation.
@@
Figure 1: Model architecture. GLiDRE uses a dual-encoder structure: a text encoder generates token embeddings, while a label encoder computes embeddings for relation types. An entity’s representation is formed by pooling the embeddings of its constituent tokens and mentions. For each entity pair, their representations are concatenated and processed by a feed-forward network before final scoring. The model can optionally incorporate a localized context pooling module, inspired by ATLOP, to enrich the relation representation.
@@
However, if recent works tackle sentence-level relation extraction in low-resource settings (Boylan et al., 2025, Li et al., 2024a, Lan et al., 2023), DocRE in similar settings remains largely unexplored. Although latest LLMs show strong capabilities in zero-shot settings for tasks such as NER and RE (Sainz et al., 2024, Zhou et al., 2023, Wang et al., 2023, Wei et al., 2023), their performance on zero-shot DocRE, especially on the DocRED or Re-DocRED dataset is still limited (Li et al., 2023, Xue et al., 2024). The recent success of GLiNER (Zaratiana et al., 2024) has motivated this work. GLiNER deals with the task of NER as matching entity type representations to textual span representations in the latent space. This approach allows better representations than LLMs due to its use of bidirectional encoders and it solves the scalability issues of autoregressive models, surpassing much larger LLMs in zero shot settings. Moreover, as a zero-shot model, GLiNER is not restricted to a fixed set of relation labels and can generalize to arbitrary relation types at inference time. Based on this work, we introduce GLiDRE, a model inspired by the bi-encoder version of GLiNER [^2] https://blog.knowledgator.com/meet-the-new-zero-shot-ner-architecture-30ffc2cb1ee0, adapted for DocRE. Our contributions in this paper are :
  • We propose GLiDRE, a new and efficient model for document-level relation extraction, inspired by the core principles of GLiNER. We also adapt and test key components from ATLOP to improve relation representations.
@@
  • We conduct extensive experiments on the Re-DocRED benchmark, demonstrating that GLiDRE achieves state-of-the-art results in challenging few-shot scenarios.
@@ 2.1 Document Level Relation Extraction
Let a document be represented as a sequence of tokens $D$. Within this document, there is a set of $m$ pre-identified entities $E=\{e_{1},e_{2},\dots,e_{m}\}$, where each entity $e_{i}$ is associated with one or more mentions in the text.
The objective of DocRE is to identify valid relational triplets. Each triplet is of the form $(e_{h},r,e_{t})$, where $e_{h},e_{t}\in E$ are the head and tail entities respectively ($h$ can be equal to $t$), and $r$ is the relation that holds between them. Based on this definition, DocRE is as a multi-label classification task.
@@ 2.2 GLiDRE architecture
GLiDRE extends the bi-encoder variant of GLiNER. This architecture employs two separate transformers, one for encoding the document and another for encoding relation labels, which offers several benefits : unlike the uni-encoder version, the bi-encoder approach avoids concatenating labels and document tokens into a single input. Instead, we encode the document and labels separately. This eliminates the need for special separation tokens and overcomes encoder context length limits, permitting an effectively unlimited label set. Moreover, label embeddings can be precomputed, accelerating inference. However, the bi-encoder architecture increases memory usage and lacks cross-attention between labels. As a result, it may struggle to disambiguate semantically similar labels.
@@
Label embeddings are obtained by mean-pooling over the token representations constituting label names. The computed label embeddings are passed through a two-layer feedforward network if the dimension of the label encoder’s output differs from the configured latent space dimension of GLiDRE. For document tokens, we adopt the standard strategy used in GLiNER and other standard NER models : for words split into subwords, we extract the representation of the first subword.
@@
First, for each mention of an entity, we compute the mention representation by pooling the embeddings of its constituent words.
@@ 2.4 Localized Context Pooling
The final relation representation $h_{r}$ is then computed using Equation 1.
@@ 2.5 Pretraining data
To construct a large-scale pre-training corpus, we use a semi-automated annotation methodology inspired by Zhou et al. (2023). We randomly sample documents from FineWeb Penedo et al. (2024), a high-quality dataset built from filtered and deduplicated English Common Crawl archives.
We prompt the Mistral-Small-24B-Instruct-2501 model to generate annotations for both entities and the relations between them in a structured JSON format (the detailed prompt is provided in appendix A.1). The raw outputs are filtered to remove instances containing malformed JSON or documents that exceed a predefined length threshold.
@@ 3 Experiments
3.1 Dataset
We conduct our experiments on the English benchmark Re-DocRED Tan et al. (2022b), a human-revised version of the DocRED benchmark that addresses its high false-negative rate, logical inconsistencies and coreference errors by re-annotating all 4,053 documents in the original training and evaluation splits. Re-DocRED retains the same relation schema and splits but substantially increases the number of annotated relation triples per document.
@@ 3.2 Evaluation Settings
We evaluate the performance of our model across three distinct experimental settings, with a particular focus on low-data regimes where its advantages are most pronounced. First, in a fully supervised setting, we fine-tune on the entire training set and compare against strong baselines. Second, to assess data efficiency and few-shot generalization, we fine-tune on randomly sampled subsets of the training data containing 1, 5, 10, 50, 100, 500, or 1000 examples. Finally, we examine zero-shot performance by comparing against larger LLMs without any task-specific fine-tuning, highlighting the model’s competitiveness in scenarios with little or no labeled data.
@@
We use the standard metrics from the Re-DocRED dataset, defined in Tan et al. (2022b): F1 refers to the micro-averaged F1 score on the recognition of relation triples, Ign_F1 corresponds to the F1 score when ignoring the triples in the test set that appear also in the training set.
@@
All experiments are conducted on a single NVIDIA H100 GPU with a batch size of 16. For the pre-training phase, the model is trained for 50,000 steps. For finetuning, we train for 10,000 steps. We use two learning rates $1\times 10^{-5}$ for the model’s encoders and $1\times 10^{-4}$ for all other layers. For the few-shot and fully supervised experiments, we report the average and standard deviation over 5 runs with different seeds. The model checkpoint yielding the best performance on the development set and a decision threshold of 0.5 are used for the evaluation on the test set. We use two English models : DebertaV3 Large He et al. (2021) for document encoder and BGE-Large V1.5 Xiao et al. (2023) for label encoder similar to the bi-encoder version of GLiNER. The final model comprises approximately 800 million parameters. The pre-training phase requires approximately 24 hours, while finetuning on the ReDocRED dataset takes 3.5 hours.
@@ 3.3 Results
3.3.1 Fully supervised setting
We evaluate the performance of our proposed model, GLiDRE, by comparing it against several strong baselines on the Re-DocRED dataset under a fully supervised protocol. Specifically, we include:
  • ATLOP (Zhou et al., 2021): Formulates document-level relation extraction as a semantic-segmentation task, introducing localized context pooling to capture entity-focused local contexts and an adaptive thresholding loss to learn dynamic decision thresholds.
  • KD-DocRE (Tan et al., 2022a): Leverages an axial attention module to model inter-entity dependencies across sentences, applies an adaptive focal loss to mitigate class imbalance and uses knowledge distillation to incorporate distantly supervised data.
@@ 3.3.1 Fully supervised setting
  • DREEAM (Ma et al., 2023): Guides the model’s attention with evidence sentences as supervisory signals in a memory-efficient manner and employs a self-training strategy to learn evidence retrieval without explicit annotations.
  • TTM-RE (Gao et al., 2024): Introduces a Token Turing Machine memory module that augments document representations with external memory tokens, paired with a noise-robust loss tailored for positive–unlabeled distantly supervised settings.
Additionally, we benchmark against GLiREL (Boylan et al., 2025), a GLiNER sentence-level relation extraction adaptation. It is important to note that GLiREL was originally conceived for sentence-level relation extraction and operates with a shorter context length. For document-level predictions, it relies on aggregating relations based on gold-standard coreference information.
We also benchmark against recent methods that leverage LLMs. Specifically, we compare with the framework introduced by Li et al. (2024b), which involves fine-tuning Llama models (Touvron et al., 2023) using Low-Rank Adaptation (Hu et al., 2022) and enhancing the model’s performance by employing a classifier to exhibit potential relations and guide the fine-tuned LLM.

### v2

@@ Abstract
Relation Extraction (RE) is a fundamental task in Natural Language Processing, and its document-level variant poses significant challenges, due to complex interactions between entities across sentences. While supervised models have achieved strong results in fully resourced settings, their behavior with limited training data remains insufficiently studied. We introduce GLiDRE, a new compact model for document-level relation extraction, designed to work efficiently in both supervised and few-shot settings. Experiments in both low-resource supervised training and few-shot meta-learning benchmarks show that our approach outperforms existing methods in data-constrained scenarios, establishing a new state-of-the-art in few-shot document-level relation extraction. Our code will be publicly available [^1] https://github.com/robinarmingaud/glidre.
@@ 1 Introduction
DocRE presents a more realistic and complex scenario, especially relevant for real-world applications such as biomedical or industrial Information Extraction, where relationships often span sentence boundaries and the entity pairs of interest are not pre-identified. The complexity of DocRE is due to the quadratic growth of negative pairs with the number of entities in a document. Evaluation is commonly conducted on datasets like DocRED or Re-DocRED (Yao et al., 2019; Tan et al., 2022b).
@@
Figure 1: Model architecture. GLiDRE uses a dual-encoder structure: a text encoder generates token embeddings, while a label encoder computes embeddings for relation types. An entity representation is formed by pooling the embeddings of its constituent tokens and mentions. For each entity pair, their representations are concatenated and processed by a feed-forward network before final scoring. The model can optionally incorporate a localized context pooling module (Zhou et al., 2021) to enrich the relation representation.
@@
However, if recent works tackle sentence-level relation extraction in low-resource settings (Boylan et al., 2025; Li et al., 2024a; Lan et al., 2023), there has been little research on DocRE in similar settings. In zero-shot settings, LLMs show strong capabilities for tasks such as NER and RE (Sainz et al., 2024; Zhou et al., 2023; Wang et al., 2023; Wei et al., 2023), but their performance on zero-shot DocRE is still limited (Li et al., 2023, Xue et al., 2024). On the other hand, lightweight zero-shot models have also achieved great performances for NER, such as GLiNER (Zaratiana et al., 2024), which exploits similarities between entity type representations and textual span representations in the latent space. This approach allows better representations than LLMs due to its use of bidirectional encoders and it solves the scalability issues of autoregressive models, surpassing much larger LLMs in zero-shot settings. Moreover, as a zero-shot model, GLiNER is not restricted to a fixed set of relation labels and can generalize to arbitrary relation types at inference time. With the same inspiration, we offer the following contributions:
  • We propose GLiDRE, a new and efficient model for document-level relation extraction with zero or few annotated data, that leverages embedding similarities between relation types and entity pair representations.
@@
  • We conduct extensive experiments on the Re-DocRED, FREDo and Re-FREDo benchmarks, demonstrating that GLiDRE achieves state-of-the-art results in challenging few-shot scenarios.
@@
2 Related Work
Document-level RE methods
State-of-the-art supervised approaches to DocRE generally extend the model ATLOP (Zhou et al., 2021), which introduces Localized Context Pooling, using attention mechanisms to enhance relation representations, and an adaptive-thresholding loss that adjusts decision thresholds dynamically based on entity pairs. Subsequent work has extended ATLOP in various directions. For example, DREEAM (Ma et al., 2023) uses evidences to enhance performance, while KD-DocRE (Tan et al., 2022a) uses axial attention to model interdependencies among entity pairs, adopts an adaptive focal loss to address class imbalance, and integrates knowledge distillation. Other works enhance DocRE models with rule constraints, adding logical constraints to improve consistency, with techniques such as bidirectional rule mining Liu et al. (2023) or differentiable document-specific rule learning Zhang et al. (2025). Another trend leverages LLMs: for instance, LMRC Li et al. (2024b) combines a dedicated classifier to suggest candidate relations with the generative ability of a LLM for final decisions, significantly improving upon purely generative approaches.
Benchmarks for DocRE
Document-level Relation Extraction models are commonly evaluated on the benchmark DocRED Zhou et al. (2021), and its refined version Re-DocRED (Tan et al., 2022b), which reduces the issue of false negatives. These two datasets contain complex documents with an average of 19 entities and approximately 390 candidate pairs per text, for a total of 3,053 annotated Wikipedia articles in the training set. Beyond these benchmarks, other datasets have been proposed for specialized domains, such as HacRED (Cheng et al., 2021) for Chinese relation extraction and SciERC (Luan et al., 2018) for scientific literature. However, to the best of our knowledge, DocRED and Re-DocRED remain the only widely adopted English benchmarks that provide evidence annotations, an essential feature for fair comparison with evidence-aware models such as DREEAM.
Few-Shot Document-level Relation Extraction
Given the complexity of document-level relation extraction, the few-shot setting has received comparatively limited attention. A straightforward strategy involves training under data-constrained regimes, where only a small number of labeled samples are available. Moving beyond this paradigm, Popovic and Färber (2022) formulate Few-Shot Document-level Relation Extraction (FSDRE) as a meta-learning problem and introduce the FREDo dataset, constructed upon DocRED and SciERC. Models are evaluated on episodic tasks comprising support and query sets derived from DocRED in the original FREDo, and from Re-DocRED in its refined version, Re-FREDo (Meng et al., 2023). SciERC is employed as an out-of-domain test set to evaluate cross-domain generalization. Approaches evaluated on these benchmarks, such as RAPL (Meng et al., 2023) and TPN (Zhang and Kang, 2024), typically employ prototype-based learning to derive robust class representations from only a handful of examples.
@@ 3.1 Document Level Relation Extraction
Let a document be represented as a sequence of tokens. Within this document, there is a set of $m$ pre-identified entities $E=\{e_{1},e_{2},\dots,e_{m}\}$, where each entity $e_{i}$ is associated with one or more mentions in the text.
The objective of DocRE is to identify valid relational triplets. Each triplet is of the form $(e_{h},r,e_{t})$, where $e_{h},e_{t}\in E$ are the head and tail entities respectively ($h$ can be equal to $t$), and $r$ is the relation that holds between them. Based on this definition, DocRE is a multi-label classification task on every entity pair, where labels are the relation types.
@@ 3.2 GLiDRE architecture
GLiDRE architecture is inspired by the bi-encoder variant of GLiNER [^2] https://blog.knowledgator.com/meet-the-new-zero-shot-ner-architecture-30ffc2cb1ee0. The GLiDRE model, presented in Figure 1, uses two separate transformers, one for encoding the document and another for encoding relation labels, which offers several benefits: unlike the uni‑encoder version, the bi-encoder approach avoids concatenating labels and document tokens into a single input. Instead, we encode the document and labels separately. This eliminates the need for special separation tokens and overcomes encoder context length limits, permitting an effectively unlimited label set. Moreover, label embeddings can be precomputed, accelerating inference. However, the bi‑encoder architecture increases memory usage and lacks cross‑attention between labels. As a result, it may struggle to disambiguate semantically similar labels.
@@
Label embeddings are obtained by mean-pooling over the token representations constituting label names. The computed label embeddings are passed through a two-layer feedforward network if the output dimension of the label encoder differs from the configured latent space dimension of GLiDRE. For document tokens, we adopt the strategy used in standard NER models: for words split into subwords, we extract the representation of the first subword.
@@
First, for each mention of an entity, the mention representation is computed by pooling the embeddings of its constituent words.
@@ 3.4 Localized Context Pooling
The final relation representation $h_{r}$ is then computed from $h^{\prime}_{e_{h}}$ and $h^{\prime}_{e_{t}}$ using Equation 1.
@@ 3.5 Pretraining data
For a better generalization, GLiDRE benefits from pretraining on a large-scale generalist annotated corpus. We build such a corpus using a semi-automated annotation methodology inspired by Zhou et al. (2023). Documents are first randomly sampled from FineWeb Penedo et al. (2024), a high-quality dataset built from filtered and deduplicated English Common Crawl archives.
We then prompt a LLM to generate annotations for both entities and the relations between them in a structured JSON format (we used the Mistral-Small-24B-Instruct-2501 model, the detailed prompt is provided in Appendix A.1). The raw outputs are filtered to remove instances containing malformed JSON or documents that exceed a predefined length threshold.
@@ 4 Experiments
4.1 Datasets
We conduct our experiments on the English benchmark Re-DocRED Tan et al. (2022b), a human-revised version of the DocRED benchmark that addresses its high false-negative rate, logical inconsistencies and coreference errors by re‑annotating all 4,053 documents in the original training and evaluation splits. Re-DocRED retains the same relation schema but substantially increases the number of annotated relation triples per document.
For evaluating performance in a formal few-shot meta-learning context, we additionally use the FReDo (Popovic and Färber, 2022) and Re-FReDo (Meng et al., 2023) benchmarks. These datasets re-purpose documents from DocRED, Re-DocRED, and SciERC into an episodic format. The evaluation is structured into two tasks: an in-domain task with approximately 15k episodes derived from DocRED, and a more challenging cross-domain task with 3k episodes derived from SciERC. Each episode consists of a small support set (1 or 3 documents) and a query set. To ensure a rigorous few-shot evaluation, the relation types in the training, development, and test splits are disjoint.
@@ 4.2 Evaluation Settings
Standard supervised settings
We evaluate the performance of our model on Re-DocRED across 3 distinct experimental settings, with a particular focus on low-data regimes where its advantages are most pronounced. First, to assess data efficiency generalization, we finetune different models on randomly sampled subsets of the training data containing 1, 5, 10, 50, 100, 500, or 1000 documents. Second, in a fully supervised setting, we finetune on the entire training set and compare against strong baselines. Finally, we examine zero-shot performance by comparing against larger LLMs without any task-specific fine-tuning, highlighting the competitiveness of the model in scenarios with little or no labeled data.
Episodic meta-learning settings
To evaluate our model in the FSDRE setting on the FReDo and Re-FReDo benchmarks, we adopted a two-stage fine-tuning protocol for a fair comparison with prototype-based methods. First, we perform an initial alignment phase by training the model for 500 steps on the full training set, mainly to align label representations (e.g., capitalization consistency). Subsequently, for each episode, we further finetune the model for 20 epochs on the provided 1-shot or 3-shot support set before evaluating its performance on the corresponding query set.
@@
We use the standard metrics for the Re-DocRED dataset (Tan et al., 2022b): F1 is the micro-averaged F1 score on the relation triples, Ign_F1 is the F1 score ignoring the triples in the test set that appear also in the training set.
@@
All experiments are conducted on a single NVIDIA H100 GPU with a batch size of 16. For the pretraining phase, the model is trained for 50,000 steps. For finetuning, we train for 10,000 steps. We use two learning rates $1\times 10^{-5}$ for the model encoders and $1\times 10^{-4}$ for all other layers. For the low-resource and fully supervised experiments, we report the average and standard deviation over 5 runs with different seeds. The model checkpoint yielding the best performance on the development set and a decision threshold of 0.5 are used for the evaluation on the test set. We use two English models: DebertaV3 Large He et al. (2021) for document encoder and BGE-Large V1.5 Xiao et al. (2023) for label encoder, similar to the bi-encoder version of GLiNER. The final model comprises approximately 800 million parameters. The pretraining phase requires approximately 24 hours, while finetuning on the ReDocRED dataset takes 3.5 hours. Additionally, evaluating our model on the Re-FReDo benchmark is computationally intensive due to the episodic protocol, which requires repeated fine-tuning and model weight resets. This process amounts to approximately 35 hours of computation for the in-domain task and 5 hours for the cross-domain task.
@@ 4.3 Results
We present in this section the main results of the evaluation of GLiDRE in several settings. An additional detailed analysis of some parameters of the model such as the effect of pretraining, pooling strategies and adaptive thresholding are presented in Appendix A.3.
4.3.1 Low-Resource Settings
In low-resource scenarios, we evaluate the efficiency of our model by comparing it against two strong supervised baselines:
@@ 4.3.1 Low-Resource Settings
DREEAM (Ma et al., 2023) represents state-of-the-art on the fully-supervised benchmark and relies on guiding the model attention with evidence information as supervisory signals in a memory‐efficient manner and can employs a self‐training strategy to learn evidence retrieval without explicit annotations.
ATLOP (Zhou et al., 2021) formulates document-level relation extraction as a semantic‐segmentation task, introducing localized context pooling to capture entity‐focused local contexts and an adaptive thresholding loss to learn dynamic decision thresholds. It is more comparable to our approach since it does not requires additional evidence information.
This comparison is designed to determine the dataset size at which these conventional supervised methods can match the performance of our model and quantifying the advantages of our approach in data-scarce regimes.
Model N = 1 N = 5 N = 10 N = 50 N = 100 N = 500 N = 1000
GLiDRE (Ours) $\mathbf{24.45}_{\pm 6.37}$ $\mathbf{33.71}_{\pm 6.11}$ $\mathbf{41.73}_{\pm 2.94}$ $\mathbf{54.54}_{\pm 1.39}$ $\mathbf{60.04}_{\pm 0.50}$ $69.26_{\pm 0.37}$ $\mathbf{72.09}_{\pm 0.31}$
Table 1: Low-resource supervised setting: micro-averaged F1 results for GLiDRE, ATLOP and DREEAM across various training set sizes (NN). All models are trained on the same 5 subsets for each value of $N$. Best results per column are shown in bold.
The results presented in Table 1 clearly demonstrate the superiority of GLiDRE in data-scarce environments. In the extremely low-data regime (N $\leq$ 100), our model establishes a lead over both ATLOP and the state-of-the-art model, DREEAM. For instance, with only 10 training samples, GLiDRE achieves an F1 score of $41.73$, surpassing DREEAM by a margin of over 14 F1 points. This significant advantage persists up to N=100 samples, the performance gap narrowing only when the amount of training data increases to 500 and 1000 samples.
4.3.2 Few-Shot Document-level Relation Extraction
Model FREDo ReFREDo
In-Domain Cross-Domain In-Domain Cross-Domain
1-Doc $F_{1}$ 3-Doc $F_{1}$ 1-Doc $F_{1}$ 3-Doc $F_{1}$ 1-Doc $F_{1}$ 3-Doc $F_{1}$ 1-Doc $F_{1}$ 3-Doc $F_{1}$
DL-Base 0.60 0.89 1.76 1.98 1.38 1.84 1.76 1.98
DL-MNAV 7.05 8.42 0.84 0.48 12.97 12.43 1.12 2.28
$\textnormal{DL-MNAV}_{SIE}$ 7.06 6.77 1.77 2.51 13.37 12.00 1.39 2.92
$\textnormal{DL-MNAV}_{SIE+SBN}$ 1.71 2.79 2.85 3.72 4.59 5.43 2.84 3.86
RAPL 8.75 10.67 3.33 5.35 15.20 16.35 3.51 5.48
TPN 9.12 8.64 3.98 4.48 15.54 15.73 4.72 5.02
GLiDRE (Ours) 13.00 15.71 10.97 11.00 26.36 28.77 10.60 10.32
Table 2: Few-shot episodic setting: results on FREDo and ReFREDo benchmarks. Reported scores are macro-averaged across relation types and obtained from Meng et al. (2023) and Zhang and Kang (2024). Best results per column are shown in bold.
For the FSDRE evaluation, we follow the episodic protocols defined by FREDo and Re-FREDo. We compare GLiDRE to three methods specifically developed for episodic few-shot DocRE:
DL-MNAV (Popovic and Färber, 2022) adapts sentence-level few-shot relation extraction MNAV Sabo et al. (2021) to documents by pooling mention representations and explicitly modeling NOTA (none-of-the-above) via learned NOTA vectors, adaptive-threshold loss, and support-based NOTA sampling at inference.
RAPL (Meng et al., 2023) refines prototypes with instance-level aggregation and relation-weighted contrastive learning, and constructs task-specific NOTA prototypes.
TPN (Zhang and Kang, 2024) improves cross-domain transfer with a hybrid encoder, transferable NOTA prototype learning, and a calibration module to mitigate NOTA bias.
As shown in Table 2, GLiDRE establishes a new state-of-the-art on both benchmarks by a substantial margin across all settings. The advantage of our model is particularly pronounced in the challenging cross-domain setting, where the task is to generalize to relations from the SciERC dataset, highlighting the effectiveness of our synthetic data pretraining strategy for FSDRE and domain transfer.